40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Feb 20 23:55:51 2019, gautam, Update, ALS, ALS delay line electronics delayLineDemod.pdf.zip
    Reply  Thu Mar 7 01:06:38 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Tue Mar 12 22:51:25 2019, gautam, Update, ALS, ALS delay line electronics 
       Reply  Wed Mar 13 01:27:30 2019, gautam, Update, ALS, ALS delay line electronics DFDcal.pdfDFDnoise.pdf
          Reply  Sun Mar 17 00:42:20 2019, gautam, Update, ALS, NF1611 cannot be shot-noise limited? PDnoise.pdf
             Reply  Thu Mar 28 19:40:02 2019, gautam, Update, ALS, BeatMouth with NF1611s assembled IMG_7381.JPGIMG_7382.JPGrelTF_schem.pdfdarkNoise.pdf
                Reply  Fri Mar 29 21:00:06 2019, gautam, Update, ALS, BeatMouth with NF1611s installed IMG_7384.JPGIMG_7385.JPG
                   Reply  Sun Mar 31 15:05:53 2019, gautam, Update, ALS, Fiber beam-splitters not PM IMG_7384.JPG
                      Reply  Wed Apr 3 09:04:01 2019, gautam, Update, ALS, Note about new fiber couplers 
                         Reply  Wed Apr 3 12:32:33 2019, Koji, Update, ALS, Note about new fiber couplers 
                      Reply  Fri Apr 5 00:33:58 2019, gautam, Update, ALS, Promising IR ALS noise IMG_7388.JPGfreeSpace_IR_beat.pdf
                         Reply  Fri Apr 5 11:49:30 2019, gautam, Update, ALS, PSL + X green beat recovery 
                         Reply  Mon Apr 8 00:04:08 2019, gautam, Update, ALS, IR ALS noise budget DFDnoise.pdf
                            Reply  Mon Apr 8 18:28:25 2019, gautam, Update, ALS, EX Green PDH checkout IMG_7393.JPGgreenModeMatch.pdf
                               Reply  Mon Apr 8 23:52:09 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated errMonCalib.pdferrMon.pngOLTF.pdfEX_frequencyNoises.pdf
                                  Reply  Tue Apr 9 00:16:22 2019, rana, Update, ALS, EX Green PDH error monitor calibrated 
                                     Reply  Tue Apr 9 00:18:19 2019, gautam, Update, ALS, EX Green PDH error monitor calibrated 
                                  Reply  Tue Apr 9 18:44:00 2019, gautam, Update, ALS, EX Green PDH discriminant measurement cavityFlashes.pdfmodelPDH.pdf
                            Reply  Tue Apr 9 19:07:12 2019, gautam, Update, ALS, IR ALS noise budget ALS_noiseBudget.pdfALS_noiseBudget.pdf
                               Reply  Thu Apr 11 01:10:05 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X EX_PDH_2kNoise.pdf
                                  Reply  Wed Apr 17 00:50:17 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDHnoise.pdf
                                     Reply  Wed Apr 17 11:01:49 2019, gautam, Update, ALS, Large 2kHz peak (and harmonics) in ALS X no more EX_PDH_specGram.pdf
                      Reply  Wed May 29 18:13:25 2019, gautam, Update, ALS, Fiber beam-splitters are now PM 
                         Reply  Fri May 31 15:55:16 2019, gautam, Update, ALS, PSL + X beat restored 
Message ID: 14521     Entry time: Mon Apr 8 00:04:08 2019     In reply to: 14516     Reply to this: 14523   14528
Author: gautam 
Type: Update 
Category: ALS 
Subject: IR ALS noise budget 

To start the noise budgeting, I decided to measure the "DFD noise", which is really the quadrature sum of the following terms:

  • ZHL-3A (RF amplifier) noise, NF ~ 6dB per spec (~ 1nV/rtHz)
  • Delay line demod board noise, ~30nV/rtHz [measurement]
  • AA board noise [measurement]
  • ADC noise

According to past characterizations of these noises, the ADC noise level, which is expected to be at the level of a few uV/rtHz, is expected to be the dominant noise source.

The measurement was made by disconnecting the NF 1611 free space photodiode from the input to the RF amplifier on the PSL table, and connecting a Marconi (f_carrier = 40 MHz, signal level=-5dBm) instead. The phase tracked output was then monitored, and the resulting digital spectrum is the red curve in Attachment #1. The blue curve is the ASD of fluctuations of the beatnote between the PSL and EX IR beams, as monitored by the DFD system, with the X arm cavity length locked to the PSL frequency via the LSC servo, and the EX green frequency locked to the X arm cavity length by the analog PDH servo. 

Conclusions:

Assuming the Marconi frquency noise is lower than the ones being budgeted:

  1. the measured frequency noise is above the DFD noise - this needs to be budgeted.
  2. The DFD noise level is consistent with a frequency discriminant of 15 uV/rtHz and an ADC noise level of 3 uV/rtHz at high frequencies.

Next noises to budget:

  1. In-loop X arm length noise
  2. In-lop EX laser frequency noise
Attachment 1: DFDnoise.pdf  62 kB  Uploaded Mon Apr 8 01:12:53 2019  | Hide | Hide all
DFDnoise.pdf
ELOG V3.1.3-